# GOFA: A Generative One-For-All Model for Joint Graph Language Modeling

### v1

@@ Abstract
Foundation models, such as Large Language Models (LLMs) or Large Vision Models (LVMs), have emerged as one of the most powerful tools in the respective fields. However, unlike text and image data, graph data do not have a definitive structure, posing great challenges to developing a Graph Foundation Model (GFM). For example, current attempts at designing general graph models either transform graph data into a language format for LLM-based prediction or still train a GNN model with LLM as an assistant. The former can handle unlimited tasks, while the latter captures graph structure much better—yet, no existing work can achieve both simultaneously. In this paper, we identify three key desirable properties of a GFM: self-supervised pretraining, fluidity in tasks, and graph awareness. To account for these properties, we extend the conventional language modeling to the graph domain and propose a novel generative graph language model GOFA to solve the problem. The model interleaves randomly initialized GNN layers into a frozen pre-trained LLM so that the semantic and structural modeling abilities are organically combined. GOFA is pre-trained on newly proposed graph-level next-word prediction, question-answering, and structural tasks to obtain the above GFM properties. The pre-trained model is further fine-tuned on downstream tasks to obtain task-solving ability. The fine-tuned model is evaluated on various downstream tasks, demonstrating a strong ability to solve structural and contextual problems in zero-shot scenarios. The code is available at https://github.com/JiaruiFeng/GOFA.
@@ 1 Introduction
With the emergence of Large Language Models (LLMs), the field of artificial intelligence is undergoing a profound transformation, shifting from specialized, fragmented models to universal foundation models. A foundation model is pre-trained on large-scale datasets and can be further adapted to diverse downstream tasks using fine-tuning Hu et al. (2022) or in-context learning Bommasani et al. (2021); Touvron et al. (2023). Foundation models have been developed in different domains to handle text Brown et al. (2020); Touvron et al. (2023), image Kirillov et al. (2023); Bai et al. (2023), and even multi-modal data Zhang et al. (2023c); Li et al. (2023); Alayrac et al. (2022). Because of their versatility and generalizability, these foundation models have become prevalent in these domains.
However, despite preliminary efforts, a foundation model in the graph domain has yet to be proposed. In the graph domain, data are highly flexible and dynamic. For example, social networks receive millions of new connections daily Hardiman and Katzir (2013), and novel molecules and protein structures are frequently discovered Abramson et al. (2024); Gilmer et al. (2017). While past researchers have proposed specialized models to learn graph data Ying et al. (2021); Kipf and Welling (2017), the models require retraining to accommodate new graphs Dai et al. (2022); Mo et al. (2022). Moreover, trained models are usually tied to specific applications and cannot be generalized to new domains and tasks. It becomes increasingly difficult for models to adjust to the ever-evolving nature of graph data. Hence, a graph foundation model that requires minimal data to adapt to new domains/tasks is urgently needed, which has spurred recent endeavors to study general graph models.
The success of LLM inspired a series of preliminary attempts to use LLM to develop general graph models. They can be roughly divided into two categories: LLM as a predictor and LLM as an enhancer Chen et al. (2023). The LLM as a predictor approach transforms graph data into representations that LLM can understand and use LLM to generate predictions Tang et al. (2023); Chen et al. (2024). However, as suggested by a recent study Wang et al. (2023), such an approach falls short of understanding graph structures. This inspired the LLM as an enhancer approach which adopts LLM to process and unify diverse graph data and feeds them to a GNN to train general graph models Liu et al. (2023a); Huang et al. (2023a). Nevertheless, because GNN outputs fixed-sized representations, they can only handle specific tasks such as classification, and cannot generalize to a foundational level due to the lack of generation ability. In summary, the current two approaches cannot fully utilize structural information and be generative simultaneously. We discuss the pros and cons of existing approaches in detail in Section 2.
@@
In this paper, we first identify three desirable properties of a graph foundation model (GFM), namely large-scale self-supervised pretraining, fluidity in tasks, and graph understanding. To achieve the first property, we propose a generic graph self-supervised learning problem similar to the next-token prediction problem in LLMs, allowing label-agnostic and continuous training on any graph data. We then propose a generative model termed Generative One-For-All (GOFA) that interleaves GNN layers into an LLM to achieve the second and third properties. Such a design painlessly integrates GNN into an LLM, granting the LLM graph structural learning ability while keeping LLM’s original free-form text generation ability. Meanwhile, this design allows the pipeline of the original LLM to remain intact, giving the GOFA a close-to-LLM level of task fluidity. We pre-train the model with large-scale real-world graph data, Question-Answer (QA) chain data adopted from the NLP domain, and graph structural data to empower the model with the aforementioned foundational abilities in the graph domain (Examples in Figure 1). After pre-training, we further fine-tune the model on downstream tasks. The fine-tuned model is evaluated on various datasets and tasks. Particularly, the GOFA achieved great results on the zero-shot scenario, which demonstrates the strong potential of the GOFA to serve as a graph foundation model.
@@ 2 A Desired Foundation Model for Graph
In this section, we summarize three crucial properties a true graph foundation model should possess to motivate our GOFA model design.
@@
Large-scale Self-Supervised Pretraining: One fundamental design of LLM is that it unifies all NLP tasks into a single next-token-prediction paradigm, which enables self-supervised pretraining on large corpus collected from different sources. For pre-training graph models, while numerous efforts have been made from both the LLM as a predictor and LLM as an enhancer approaches, these attempts usually require the learning target to be labeled Liu et al. (2023a); Chen et al. (2023). However, a graph foundation model should have no constraint on the input graph (has labels or not) and can learn cross-domain knowledge from large-scale graph data in an unsupervised/self-supervised fashion.
@@
Fluidity in Tasks: A graph foundation model should also possess the same level of versatility and fluidity in handling different tasks as an LLM. Specifically, such ability can be broken down into three levels: (a) The graph foundation model can naturally respond appropriately to different trained graph tasks according to the user instructions. (b) With appropriate instruction-tuning, the model should have in-context learning ability on unseen tasks. (c) Users can define new tasks unseen in training data by modifying the graph structure and features that fit into the universal input representation of the model. They can continuously train the model on new data without special adaptation. Existing approaches that use GNN models as the predictors are usually either restricted in the output format Liu et al. (2023a); Xia et al. (2024); He et al. (2024a) or need additional fine-tuning on the task head Sun et al. (2023); Wang et al. (2022). Consequently, despite having better structural modeling ability, such models cannot accommodate task changes or deal with novel tasks, e.g., shifting from a classification task to a regression task or to a question-answering task that requires outputting all shortest paths between two nodes.
@@
Graph Understanding: Since the LLM as a predictor approach uses a generative LLM to produce output, it naturally has the fluidity to accept varied prompts to tackle different tasks. However, such an approach processes the structural information poorly Wang et al. (2023), making the utility of these models minimal on graph. More importantly, even though some recent variants can use auxiliary graph models (such as GNNs) to incorporate structural information Tang et al. (2023); He and Hooi (2024); Zhang et al. (2024), the graph models are frozen and not responsive to different prompts, and the output from the graph models may not be the most relevant to the input prompt. On the contrary, a graph foundation model should account for the unique structural information of graphs such as node degrees, shortest paths, common neighbors, etc., and generate graph representations dependent on the input prompt. It should not only have LLM’s prompt learning capability but also learn graph structure and semantic information jointly.
@@ 3 Method
In this section, we first propose a generative modeling framework for graph, serving as the graph counterpart of the traditional language modeling. Next, we introduce a novel GNN-LLM architecture for the proposed graph generative modeling problem. Finally, we describe the unified pre-training and downstream fine-tuning process of the model.
@@ 3.1 Generative Modeling for Graph
In plain text applications, aligning the pre-training objective with the input and output format of downstream tasks is straightforward, as they are purely text-based. However, graph data from different domains vary significantly by input feature dimensions and output target. Hence, the first challenge is to define a unified input and output format for graph tasks, such that the model can do large-scale self-supervised pre-training on arbitrary graphs and downstream tasks.
For the unified input, following the previous work OFA Liu et al. (2023a), we extend the definition of Text-Attribute Graph (TAG) beyond graphs that already have text features such as citation and product networks. Specifically, we observe that any node and edge features can be represented by texts. For example, atom and bond descriptions replace the node and edge features in a molecular graph. Furthermore, we can represent non-textual features like pure numbers as texts too, just like they are represented as text strings in LLM. Formally, a TAG is a graph $G=\{V,E,X_{V},X_{E}\}$ where $V$ and $E$ are the sets of nodes and edges. Each node $v\in V$ (edge $e\in E$) corresponds to a text description $x(v)\in X_{V}$ ($x(e)\in X_{E}$). Such a format encodes almost all existing graph data and serves well as the universal input representation. For the unified output, we believe that natural language is still the most tangible format for users. Such input and output formats cover almost all commonly seen graphs-related tasks, including classification, regression, and free-form question answering.
Next, we define the generative modeling framework for graphs. Generative modeling for language is to predict the next token given its preceding tokens. However, in graph-based tasks, input sentences are distributed to different nodes and edges in a TAG. To generalize next token prediction to graphs, we propose to specify Nodes Of Generation (NOG) on graphs as starting points for output generation. Formally, we can define graph generative modeling as the likelihood of the text $y$ associated with the NOG $v$:
@@
Figure 2: Task examples in TAG. Left: Sentence completion/Next-word prediction. Right: Question Answering. Node $v$ represents NOG.
where $y_{l}$ is the $l$-th token of $y$, and $y_{<l}$ is its preceding tokens. The NOG $v$ is a learning target with initial corresponding text $x(v)$, and $x(v)$ can be empty. $G$ contains structural and textual information of neighbor nodes to help the model generate $y$. For example, the sentence completion task is shown on the left of Figure 2, where the text on node $v$ is incomplete, and the goal is to complete the sentence using the existing text and the neighbor information. The question answering task is shown on the right of Figure 2, where a user’s question is injected as a node in the graph. The goal is to generate the correct answer. Note that our model is not limited to node-level tasks in conventional graph learning. We discuss more details about data construction in Section 3.3. Users can also specify multiple generation starting points for the same graph. In addition, when the graph has only one node, the problem degenerates to conventional language modeling.
@@ 3.2 GOFA : Generative One-For-All Model
To solve the generative graph modeling problem proposed in Equation 1, we design the GOFA architecture, with an illustration shown in Figure 3. GOFA consists of a graph language encoder and an LLM decoder. In the graph language encoder, we interleave token-level GNN layers with frozen LLM compressor layers. The LLM compressor independently processes the graph’s node/edge text features and outputs their compressed fixed-size multi-token embeddings which are fed into a token-level GNN layer. The GNN layer updates the multi-token embeddings by incorporating the input graph structure and feeds the output into the next LLM layer. We alternate GNN and LLM layers to get the final multi-token node representations containing joint structural and semantic information. We then use an LLM decoder to generate texts from the NOG representation. The LLM compressor and decoders are all pre-trained decoder-only transformers. We describe each component in detail as follows.
Figure 3: GOFA Framework. Text tokens of TAG’s node/edges are concatenated with memory tokens to be input to Graph Language Encoder. GNN layers are interleaved into LLM Compressor layers, where memory embeddings from LLM Compressor Layer are used as node/edge features for token-level GNN message passing. Memory embedding will be used for teacher-forcing training.
LLM compressor: Because GNNs require node and edge representations to have the same input dimension, many previous works propose to pool all tokens’ output embeddings from the LLM as the node and edge vector representations and feed them to a GNN Liu et al. (2023a); Huang et al. (2023a); He and Hooi (2024). While this approach shows effectiveness in tasks of fixed form, such as classification and regression, it is insufficient in more complex tasks, such as generation tasks, as 1) the pooling process inevitably loses semantic information, 2) standard LLMs are not trained in a way such that the pooled output embedding is a summarization of the input sentence, and 3) the pooled representation space is no longer compatible with the space of the downstream LLM decoder. Hence, we adopt a sentence compressor Ge et al. (2023) that preserves as much information as possible from the original sentence in fixed-size multi-token embeddings. Specifically, the sentence compressor layer has the same architecture as a decoder-only LLM, but the initial sequences of token representations $\{q(x_{i})\}_{i=1}^{l}$ are appended by a sequence of $K$ memory tokens $\{q(m_{j})\}_{j=1}^{K}$, and the $t$-th layer of the LLM becomes:
@@
We use $Q_{x}^{t}$ and $Q_{m,x}^{t}$ to represent the $t$-th LLM layer output corresponding to actual text tokens in $x$ and the $K$ memory tokens appended at the end of text tokens. We use $H_{x}^{t}$ to represent the $t$-th GNN layer output, which will be explained later. In Equation 2, the text tokens ($Q_{x}^{t}$) and memory tokens ($H_{x}^{t}$, processed by the previous GNN layer) are concatenated as a single sequence of embeddings, which are fed to the current LLM layer. Because the last K tokens attend to all previous tokens, they can compress all information in the sentence into the output embeddings of the K tokens. This compressor architecture is inspired by ICAE Ge et al. (2023). While the compression capability is not apparent, it is viable through auto-encoder-style fine-tuning, as discussed in Appendix A.1.
Token-level GNN: Conventional GNNs only pass messages between node vector representations, which is incompatible with the vector sequence output $Q_{m,x}$ from our LLM. Hence, we propose a simple extension of conventional GNNs to token level. For node $v\in V$, the $t$-th GNN layer is
@@
In the GNN layer, tokens at different indices do not communicate. If we directly stack these GNN layers, they degenerate into multiple isolated GNNs for each token. Nevertheless, we interleave the GNN layers into the LLM layers such that the subsequent self-attention layers on $H^{t}$ permit information flows between tokens. This approach significantly reduces memory usage because we do not perform cross-token attention in GNN. While edge memory tokens are passed into GNN to assist message passing, their representations are not updated in the GNN layer but directly passed to the next LLM layer, and $H^{t}_{x(e)}=Q^{t}_{m,x(e)}$. The specific choice of the GNN can vary. In GOFA , we use a modified Transformer Convolutional GNN Shi et al. (2021) to be consistent with the transformer architecture of LLM (see Appendix A.3 for details).
@@
We insert one GNN layer between two transformer layers, while the first and the last layer are always transformer layers. In GOFA , we only insert GNN to the last few transformer layers, but this can be flexible depending on the computational resources. Following previous practice, we incorporate feed-forward (FF) layers into the GNN to increase expressivity and residual connections to stabilize training. Inspired by the gating mechanism in earlier works Hochreiter and Schmidhuber (1997); Alayrac et al. (2022), we apply a $tanh$ gate, initialized at 0, to the GNN and FF layer outputs so that the initial model ignores the information from GNN layers and is equivalent to the pre-trained LLM.
@@
LLM decoder: After applying the model to the textual graph, the memory tokens $Q_{m,x}$ of every node contain information about the text on the node, the surrounding node text, and the graph structure due to the message-passing process in the GNN layer. Then, for the NOG $v$ and its corresponding target text $y$, we insert $Q_{m,x}$ at the front of the token embeddings of the target text to generate and use teacher-forcing to maximize the standard log-likelihood for the next-token-prediction objective.
$\mathcal{L}=\max_{Q_{m,x(v)}}P(y_{1}...y_{l}|Q_{m,x(v)};\Theta_{\text{DEC}})=\max_{\Theta_{(\text{GNN,DEC,COMP})}}P(y_{1}...y_{l}|v;G;\Theta_{\text{DEC}},\Theta_{\text{COMP}},\Theta_{\text{GNN}}).$ (4)
In this way, we have modeled the problem in Equation 1. The compressor, decoder, and GNN parameters can be jointly or separately optimized, potentially with parameter-efficient fine-tuning methods like LoRA Hu et al. (2022). In this paper, we use ICAE Ge et al. (2023) as our backbone LLM, but the GOFA architecture is not tied to any specific LLM. More details are discussed in Appendix A.2.

### v2

@@ Abstract
Foundation models, such as Large Language Models (LLMs) or Large Vision Models (LVMs), have emerged as one of the most powerful tools in the respective fields. However, unlike text and image data, graph data do not have a definitive structure, posing great challenges to developing a Graph Foundation Model (GFM). For example, current attempts at designing general graph models either transform graph data into a language format for LLM-based prediction or still train a GNN model with LLM as an assistant. The former can handle unlimited tasks, while the latter captures graph structure much better—yet, no existing work can achieve both simultaneously. In this paper, we first identify three key desirable properties of a GFM: self-supervised pretraining, fluidity in tasks, and graph awareness. To account for these properties, we extend the conventional language modeling to the graph domain and propose a novel generative graph language model GOFA. The model interleaves randomly initialized GNN layers into a frozen pre-trained LLM so that the semantic and structural modeling abilities are organically combined. GOFA is pre-trained on newly proposed graph-level next-word prediction, question-answering, structural understanding, and information retrieval tasks to obtain the above GFM properties. The pre-trained model is further instruction fine-tuned to obtain the task-solving ability. Our GOFA model is evaluated on various downstream datasets unseen during the pre-training and fine-tuning phases, demonstrating a strong ability to solve structural and contextual problems in zero-shot scenarios. The code is available at https://github.com/JiaruiFeng/GOFA.
@@ 1 Introduction
With the emergence of Large Language Models (LLMs), the field of artificial intelligence is undergoing a profound transformation, shifting from specialized, fragmented models to universal foundation models. A foundation model is pre-trained on large-scale datasets and can be further adapted to diverse downstream tasks using fine-tuning (Hu et al., 2022) or in-context learning (Bommasani et al., 2021; Touvron et al., 2023). Foundation models have been developed in different domains to handle text (Brown et al., 2020; Touvron et al., 2023), image (Kirillov et al., 2023; Bai et al., 2023), and even multi-modal data (Zhang et al., 2023c; Li et al., 2023; Alayrac et al., 2022). Because of their versatility and generalizability, foundation models have become prevalent in these domains.
However, despite preliminary efforts, a foundation model in the graph domain has arguably yet to be proposed. In the graph domain, data are highly flexible and dynamic. For example, social networks receive millions of new connections daily (Hardiman & Katzir, 2013), and novel molecules and protein structures are frequently discovered (Abramson et al., 2024; Gilmer et al., 2017). While past researchers have proposed specialized models to learn graph data (Ying et al., 2021; Kipf & Welling, 2017), the models require retraining to accommodate new graphs (Dai et al., 2022; Mo et al., 2022). Moreover, trained models are usually tied to specific applications and cannot be generalized to new domains and tasks. It becomes increasingly difficult for models to adjust to the ever-evolving nature of graph data. Hence, a graph foundation model (GFM) applicable to new domains/tasks with minimal or no adaptation costs is urgently needed, spurring recent endeavors to study general graph models. In particular, a strong zero-shot ability is both challenging and fascinating for GFM researchers.
The success of LLMs inspired a series of preliminary attempts which use LLMs to develop general graph models. They can be roughly divided into two categories: LLM as a predictor and LLM as an enhancer (Chen et al., 2023). The LLM as a predictor approach transforms graph data into representations that LLMs can understand and use LLMs to generate predictions (Tang et al., 2023). However, as suggested by a recent study (Wang et al., 2023), such an approach falls short of understanding graph structures. This inspired the LLM as an enhancer approach, which adopts LLM to process and unify diverse graph data and feeds them to a GNN to train general graph models (Liu et al., 2023a; Huang et al., 2023a). Nevertheless, because GNN outputs fixed-sized representations/predictions, they can only handle specific tasks such as classification, and cannot generalize to arbitrary, new tasks due to the lack of generation ability. In summary, the current two approaches cannot fully utilize structural information and be generative simultaneously. We discuss the pros and cons of existing approaches in detail in Section 2.
@@
In this paper, we first identify three desirable properties of a graph foundation model (GFM), namely large-scale self-supervised pre-training, fluidity in tasks, and graph understanding. To achieve the first property, we propose a generic graph self-supervised learning problem similar to the next-token prediction problem in LLMs, allowing label-agnostic and continual training on highly diverse graph data. We then propose a generative model termed Generative One-For-All (GOFA) that interleaves GNN layers into an LLM to achieve the second and third properties. Such a novel design systematically integrates GNN into an LLM, granting the LLM graph structural learning ability while keeping LLM’s original free-form text generation ability. Meanwhile, this design allows the pipeline of the original LLM to remain intact, giving GOFA a close-to-LLM level of task fluidity. We pre-train the model with large-scale real-world graph data, Question-Answer (QA) chain data adopted from the NLP domain, and graph structural data to empower the model with the aforementioned foundational abilities in the graph domain (Examples in Figure 1). After pre-training, we further instruction fine-tune the model on a small amount of data (relative to the pre-training data) to make it understand task formats. The fine-tuned model is finally evaluated on various downstream datasets unseen during pre-training and fine-tuning. GOFA achieved impressive results on the zero-shot scenario, which demonstrates the strong potential of GOFA to serve as a graph foundation model.
@@ 2 A Desired Foundation Model for Graph
In this section, we elaborate on three crucial properties a true graph foundation model should possess to motivate our GOFA model design. We note that many contemporary works (partly) propose similar ideas to ours and thus we do not claim the credit. We kindly refer readers to the latest surveys (Liu et al., 2023b; Jin et al., 2023; Zhang et al., 2023d) for more discussions on GFMs.
@@
Large-Scale Self-Supervised Pre-training: One fundamental design of LLM is that it unifies all NLP tasks into a single next-token-prediction paradigm, which enables self-supervised pre-training on a large corpus collected from different sources. For pre-training graph models, while numerous efforts have been made from both the LLM as a predictor and LLM as an enhancer approaches, these attempts usually require the learning target to be labeled (Liu et al., 2023a; Chen et al., 2023). However, a graph foundation model should have no constraint on the input graph (has labels or not) and can learn cross-domain knowledge from large-scale graph data in a self-supervised fashion.
@@
Fluidity in Tasks: A graph foundation model should also possess the same level of versatility and fluidity in handling different tasks as an LLM. Specifically, such ability can be broken down into three levels: (a) The graph foundation model can naturally respond appropriately to different graph tasks based on user instructions without requiring task-specific adjustment (e.g., the same model performs classification and question-answering tasks without any modification.) (b) With appropriate instruction-tuning, the model should have in-context learning ability on unseen tasks (e.g., a model tuned on citation network also performs well on knowledge graphs with proper instructions). (c) Users should be able to define new, previously unseen tasks by modifying the graph structure and features in a way that aligns with the universal input representation of the model. They can continuously train the model on new data without special adaptation. Existing approaches that use GNN models as the predictors are usually either restricted in the output format (Liu et al., 2023a; Xia et al., 2024; He et al., 2024a) or need additional fine-tuning on the task head (Sun et al., 2023; Wang et al., 2022). Consequently, despite having better structural modeling ability, such models cannot accommodate task changes or deal with novel tasks, e.g., shifting from a classification task to a question-answering task that requires outputting all shortest paths between two nodes.
@@
Graph Understanding: Since the LLM as a predictor approach uses a generative LLM to take text input and produce text output, it naturally has the fluidity to accept varied prompts to tackle different tasks. However, such an approach processes the structural information poorly (Wang et al., 2023), making the utility of these models limited on many graph tasks. More importantly, even though some recent variants can use auxiliary graph models (such as GNNs) to incorporate structural information (Tang et al., 2023; He & Hooi, 2024; Zhang et al., 2024), the graph models are frozen and not responsive to different prompts, and the output from the graph models may not be the most relevant to the input prompt. On the contrary, a graph foundation model should account for the unique structural information of graphs such as node degrees, shortest paths, common neighbors, etc., and generate graph representations dependent on the input prompt. It should not only have LLM’s prompt learning capability but also learn graph structure and semantic information jointly.
@@ 3 Method
In this section, we first propose a generative modeling framework for graphs, serving as the graph counterpart of traditional language modeling. Next, we introduce a novel GNN-LLM architecture for the proposed graph generative modeling problem. Finally, we describe the unified pre-training tasks to train GOFA towards the proposed GFM properties.
@@ 3.1 Generative Modeling for Graph
Unifed task formats. A generative model usually takes existing contexts, such as user prompts and passages, as input to generate conditional output related to the contexts, such as answers and completed sentences. Defining unified input and output formats for tasks in language applications is easy, as they are purely text-based. Further, because both the pre-training and downstream tasks are constructed in the same format (i.e., next-token-prediction), the downstream tasks conveniently adapt the knowledge from pre-training tasks, resulting in surprising capabilities, such as zero-shot learning. However, graph data from different domains vary significantly by input feature (e.g., nodes in a citation network have completely different vector representations as nodes in a knowledge graph) and output target, preventing direct knowledge transfer between tasks. Hence, the first challenge is to define a unified format for graph tasks, such that the model can do large-scale self-supervised pre-training on arbitrary graphs and transfer to downstream tasks seamlessly.
To unify graph task input, we follow the previous work OFA (Liu et al., 2023a) and extend the definition of Text-Attribute Graph (TAG) beyond graphs with text features such as citation and product networks. In fact, any node and edge features can be represented by texts. For example, in airline networks, airport and flight route details can be converted into textual descriptions for nodes and edges. Non-textural features, like numerical data, can also be transformed into text strings, as in LLMs. Even for graphs without any features, we can still attach sentences like "The degree of this node is 3" to nodes. Formally, a TAG is a graph $G=\{V,E,X_{V},X_{E}\}$ where $V$ and $E$ are the sets of nodes and edges. Each node $v\in V$ (edge $e\in E$) corresponds to a text description $x(v)\in X_{V}$ ($x(e)\in X_{E}$). Such a format encodes almost all existing graph data and serves well as a general input representation.
For self-supervised language modeling, the generated output essentially completes the input sentence. Such a task requires the model to have a deep semantic and logical understanding of the provided contexts, which is crucial for downstream applications. Similarly, in graph modeling, we aim to achieve the same level of understanding through graph completion tasks. Given a TAG, the output should complete the graph conditioned on its semantic and structural information. We choose to use natural language as the most tangible output format to complete a TAG. Succinctly, all natural language tasks can be modeled as sentence completion, and similarly, we aim to model all graph tasks with graph completion.
Generative Graph Modeling. We then formally define the generative graph modeling framework for graph completion. This framework supports various graph-related tasks, including classification and free-form question answering. An LLM starts generating only from the end of the input sentence. However, in a TAG, every end of a sentence on a node is a potential generation starting point, but users might only be interested in generating output for specific nodes. To accommodate this, we introduce Nodes of Generation (NOG), allowing users to specify starting points for generation. The modeling task is to take a TAG as input and complete the TAG logically and sensibly by completing the sentences on the potentially user-specified nodes.
We define graph generative modeling as the likelihood of the text $y$ associated with the NOG $v$:
@@
Figure 2: Task examples in TAG. Sentence completion/Next-word prediction. Orange node $v$ represents NOG.
where $y_{l}$ is the $l$-th token of $y$, and $y_{<l}$ is its preceding tokens. The NOG $v$ is a completion target node with initial corresponding text $x(v)$, and $x(v)$ can be empty. $G$ contains structural and textual information of neighbor nodes to help the model generate $y$. Under this framework, we can design a range of self-supervised learning tasks. For example, the graph completion task is shown on Figure 2, where the text on the NOG $v$ is incomplete, and the goal is to complete the sentence on it using the existing text and the neighbor information. This task is covered by Equation (1), which encourages the model to have a strong graph structure and feature comprehension ability. Thus, the importance of the framework is that a model properly solves such modeling problems can possess the three properties of GFM discussed in Section 2, thus can benefit diverse downstream tasks, even in the zero-shot fashion. Section F.2 discusses how the proposed framework applies to various tasks related to the three properties.
@@ 3.2 GOFA : Generative One-For-All Model
To solve the generative graph modeling problem proposed in Equation (1), we design the GOFA architecture shown in Figure 3. Overall, GOFA consists of a graph language encoder and an LLM decoder. The graph language encoder interleaves GNN layers with LLM compressor layers to learn node representations containing joint structural and semantic information. The LLM decoder is then used to generate texts from the NOG representation. The LLM compressor and decoders are all pre-trained decoder-only transformers. We describe each component in detail as follows.
Figure 3: GOFA Architecture. Text tokens of TAG’s node/edges are concatenated with memory tokens to be input to Graph Language Encoder. GNN layers are interleaved into LLM Compressor layers, where memory embeddings from LLM Compressor Layer are used as node/edge features for token-level GNN message passing. Memory embedding will be used for teacher-forcing training.
LLM compressor: Because GNNs require node and edge representations to have the same input dimension, many previous works propose to pool all tokens’ output embeddings from the LLM as the node and edge vector representations and feed them to a GNN (Liu et al., 2023a; Huang et al., 2023a; He & Hooi, 2024). While this approach shows effectiveness in tasks of fixed form, such as classification and regression, it is insufficient in more complex tasks such as generation, as 1) the pooling process inevitably loses semantic information, 2) standard LLMs are not trained in a way such that the pooled output embedding is a summarization of the input sentence, and 3) the pooled representation space is no longer compatible with the space of the downstream LLM decoder. Hence, we adopt a pre-trained sentence compressor (Ge et al., 2023) that preserves as much information as possible from the original sentence in fixed-size multi-token embeddings. The core idea is to compress a sentence into $K$ embeddings instead of one embedding. Specifically, the sentence compressor has the same architecture as a decoder-only LLM, but the sentence to be compressed $\{q(x_{i})\}_{i=1}^{l}$ is appended by a sequence of $K$ memory tokens $\{q(m_{j})\}_{j=1}^{K}$, and the $t$-th layer of the LLM is:
@@
We use $Q_{x}^{t}$ and $Q_{m,x}^{t}$ to represent the $t$-th LLM layer outputs corresponding to actual text tokens in sentence $x$ and the $K$ memory tokens appended at the end of text tokens, respectively. We use $H_{x}^{t}$ to represent the $t$-th GNN layer output, which will be explained later. In Equation (2), the text tokens ($Q_{x}^{t}$) and memory tokens ($H_{x}^{t}$, processed by the previous GNN layer) are concatenated as a single sequence of embeddings, which are fed to the current LLM layer. Because the last K tokens attend to all previous tokens, they can compress all information in the sentence into the output embeddings of the K tokens. This compressor architecture is inspired by ICAE (Ge et al., 2023). The compression ability is obtained through auto-encoder-style fine-tuning, as discussed in Appendix A.1.
Token-level GNN: Conventional GNNs take one embedding vector for each node/edge. However, now each node/edge sentence is compressed into $K$ memory token embeddings $Q_{m,x}$. Hence, we propose a simple extension of GNNs to the token level. For node $v\in V$, the $t$-th GNN layer is
@@
In the GNN layer, tokens at different indices do not communicate. If we directly stack these GNN layers, they degenerate into multiple isolated GNNs for each token. Nevertheless, because we interleave the GNN layers into the LLM layers, as shown in Figure 3, the isolated tokens exchange information in the subsequent self-attention layers of the LLM. This approach significantly reduces memory usage because we do not allow cross-token attention between different nodes. While edge memory tokens $Q^{t}_{m,x(e)}$ are passed into GNN to assist message passing, their representations are not updated in the GNN layer but directly passed to the next LLM layer, hence $H^{t}_{x(e)}=Q^{t}_{m,x(e)}$. In GOFA, we use a modified Transformer Convolutional GNN (Shi et al., 2021) to be consistent with the transformer architecture of LLM (see Appendix A.3 for details).
@@
We insert one GNN layer between two transformer layers, while the first and the last layer are always transformer layers. In GOFA, we only insert GNN between the last few transformer layers, but this can be flexible depending on the computational resources. Following previous practice, we incorporate feed-forward (FF) layers into the GNN to increase expressivity and residual connections to stabilize training. Moreover, GOFA should maintain the functions of an LLM on plain texts, hence, inspired by the gating mechanism in earlier works (Hochreiter & Schmidhuber, 1997; Alayrac et al., 2022), we apply a $tanh$ gate, initialized at 0, to the GNN and FF layer outputs so that the initial model ignores the information from GNN layers and is equivalent to the pre-trained LLM. We introduce weight decay in the gating module to promote gate value staying in large non-zero values only when graph information helps generate more accurate final text outputs.
@@
LLM decoder: After applying the model to the textual graph, the memory tokens $Q_{m,x}$ of every node contain information about the text on the node, the surrounding node text, and the graph structure due to the message-passing process in the GNN layers. Then, for the NOG $v$ and its corresponding target text $y$, we insert $Q_{m,x}$ at the front of the token embeddings of the target text to generate and use teacher-forcing to maximize the standard log-likelihood of $y$ using the next-token-prediction objective. In this way, we have modeled the problem in Equation (1). The compressor, decoder, and GNN parameters can be jointly or separately optimized, potentially with PEFT methods like LoRA (Hu et al., 2022). In this paper, we use ICAE (Ge et al., 2023) as our backbone LLM, but the GOFA architecture is not tied to any specific LLM. More details are discussed in Appendix A.2.
Discussion. Our proposed graph language encoder has several advantages over existing methods. Suppose a graph has $V$ nodes, $E$ edges, and the average number of tokens for all nodes is $k$. The complexity of one GOFA layer is $O(Vk^{2})$, as the self-attention only happens within each node. Note that we have omitted the extra computation complexity of message-passing because it only happens at individual indices with $O(E)\ll O(Vk^{2})$ in practical graphs. Instead, if we concatenate texts in all nodes and input them to a regular LLM, the complexity of one layer is $O((Vk)^{2})$, which is significantly larger than GOFA. Further, introducing GNN layers in LLMs is theoretically more powerful than pure LLMs for modeling graph structures, which is discussed in Appendix E.2.